OpenStudy (anonymous):

The height of a cylinder with a fixed radius of 6 cm is increasing at the rate of 3 cm/min. Find the rate of change of the volume of the cylinder (with respect to time) when the height is 20cm. 36π 108π 360π None of these

OpenStudy (anonymous):

@ganeshie8

OpenStudy (matlee):

C

OpenStudy (xapproachesinfinity):

First right down your cylinder volume formula \(\huge \rm \color{blue}{v=\pi r^2h}\) the radius here is fixed so when diffrentiating we get \(\huge \rm d/dx(v)=d/dx(\pi r^2h)=\pi r^2d/dx(h)\)

OpenStudy (xapproachesinfinity):

we are looking for d/dx(V) or otherwise written as V' it should d/dt not d/dx my bad

OpenStudy (xapproachesinfinity):

just plug in all the numbers you have and you are done